Find four consecutive odd integers whose sum is 416.
Answer by mananth(16949)

let the numbers be x,x+2,x+4,x+6
sum = 416
x+x+2+x+4+x+6= 416
4x+12=416
add -12 to both sides
4x+12-12=416-12
4x=404
divide by 4
4x/4 = 404/4
x= 101
the numbers are 101,103,105,107